# Self Pre-training with Topology- and Spatiality-aware Masked Autoencoders for 3D Medical Image Segmentation

### v2

@@ 4.1 Datasets and Experimental Setup
We conduct experiments on five segmentation datasets: Synapse multi-organ CT segmentation [17], BTCV multi-organ CT segmentation [17], ACDC automated cardiac diagnosis [68], and Medical Segmentation Decathlon (MSD) datasets [18] for two different segmentation tasks, spleen segmentation and lung segmentation. For each experiment, we perform 5 runs using different random seeds and report the average results. Additionally, we compute $p$-values to ascertain the statistical significance of the results.
@@
The Synapse Multi-organ CT Segmentation Dataset (Synapse CT Dataset) [^1] https://www.synapse.org/#!Synapse:syn3193805/wiki/89480: This dataset [17] contains $30$ abdominal CT volumes with $8$ organs. Following [56], we split the dataset randomly into $18$ volumes and $12$ volumes for training and testing, and report the average Dice and $95\%$ Hausdorff distance (HD95) on $8$ abdominal organs: spleen (Spl), right kidney (RKid), left kidney (LKid), gallbladder (Gal), liver (Liv), stomach (Sto), aorta (Aor), and pancreas (Pan).
@@
The BTCV Multi-organ CT Segmentation Dataset (BTCV CT Dataset)1: This dataset [17] consists of $30$ abdominal CT volumes with $13$ organs, including 8 organs of the Synapse CT dataset, along with esophagus (Eso), inferior vena cava (IVC), portal and splenic veins (PSV), right adrenal gland (RAG), and left adrenal gland (LAG). We report the average Dice scores on all the 13 abdominal organs, following [3].
@@
ACDC Automated Cardiac Diagnosis (ACDC Dataset) [^2] https://www.creatis.insa-lyon.fr/Challenge/acdc/databases.html: This dataset [68] contains 100 samples, and aims to segment the cavity of the right ventricle, the myocardium of the left ventricle, and the cavity of the left ventricle. Each sample’s labels involve left ventricle (LV), right ventricle (RV), and myocardium (MYO). Following [2], we split the dataset into 70 training samples, 10 validation samples, and 20 test samples, and use the Dice score as the evaluation metric.
@@
Spleen Segmentation on the MSD Dataset (MSD Spleen Dataset) [^3] http://medicaldecathlon.com/dataaws/: This dataset [18] contains $41$ CT volumes for spleen segmentation. Following [57], we split the dataset into training, validation, and test sets with a ratio of 80:15:5, and use the Dice score and HD95 as the evaluation metrics.
@@
Lung Segmentation on the MSD Dataset (MSD Lung Dataset)3: This dataset [18] comprises 64 CT volumes, aiming to segment lung cancer from the background. We split the dataset with a 80:20 ratio for training and testing, and use the Dice score as the evaluation metric, as in [3].
4.2 Implementation Details
@@
Our experiments are implemented with PyTorch and MONAI. [^4] https://github.com/Project-MONAI The model training is performed on an NVIDIA Tesla V100 Graphics Card with 32GB GPU memory using the AdamW optimizer [69] with a weight decay = $0.005$.
@@
For the Synapse CT and BTCV CT datasets, we clip the raw values between $-175$ and $250$, normalize the values into the range of [0, 1], and re-sample the spacing to $[1.5,1.5,2.0]$. All the models are trained with input images of size $96\times 96\times 96$. For the ACDC dataset, we re-sample the spacing to $[1.52,1.52,6.35]$. All the models are trained with input of size $160\times 160\times 16$. For the MSD spleen dataset, we clip the raw values between $-57$ and $164$, normalize the values into the range of [0, 1], and re-sample the spacing to $[1.5,1.5,2.0]$. All the models are trained with input of size $96\times 96\times 96$. For the MSD lung dataset, we clip the raw values between $-1000$ and $3071$, normalize the values into the range of [0, 1], and re-sample the spacing to $[1.0,1.0,1.0]$. All the models are trained with input of size $192\times 192\times 32$.
@@
We use a learning rate of $36.4e-$ 3 for self pre-training on all the datasets. We pre-train on the Synapse CT, BTCV CT, and MSD spleen and lung segmentation datasets with $10,000$ epochs, and on the ACDC dataset with $2,000$ epochs.
@@
For all the downstream segmentation tasks, we use a learning rate of $11e-$ 1, and fine-tune with $5,000$ epochs for the Synapse CT, BTCV CT, and MSD spleen and lung segmentation datasets, and $1,000$ epochs for the ACDC dataset. The batch size for each case is set as the maximum size allowed by the GPU. We set $\lambda_{1}=0.5$, $\lambda_{2}=0.1$, and $\lambda_{3}=0.1$.
Table 2: Segmentation results of different methods on the BTCV CT dataset. The results of the baselines are from [3]. The reported values are the averages of 5 runs with different random seeds. The best results are marked in bold, and the second-best results are underlined. We calculate the $p$-values between the average performance of our method and the best performing baseline in the average Dice score metric.
@@
[…]
$p$-value $<1e-2$ (Dice)
Table 3: Segmentation results of different methods on the ACDC dataset. The results of the baselines are from [3, 2]. ‘—” indicates that the results were not reported in the original papers, and ($\cdot$) specifies the results of our implementation of the known method. The reported values are the averages of 5 runs with different random seeds. The best results are marked in bold, and the second-best results are underlined. We calculate the $p$-values between the average performance of our method and the best performing baseline in the Dice score metric.
@@
$p$-value $<1e-2$ (Dice)
Table 4: Segmentation results of different methods on the MSD spleen dataset. The results of the baselines are from [57]. ‘—” indicates that the results were not reported in the original papers, and ($\cdot$) specifies the results of our implementation of the known method. The reported values are the averages of 5 runs with different random seeds. The best results are marked in bold, and the second-best results are underlined. We calculate the $p$-values between the average performance of our method and the best performing baseline in both the metrics.
@@
[…]
$p$-values $<1e-2$ (Dice), $<5e-2$ (HD95)
Table 5: Segmentation results of different methods on the MSD lung dataset. The results of the baselines are from [3]. “—” indicates that the results were not reported in the original papers, and ($\cdot$) specifies the results of our implementation of the known method.The reported values are the averages of 5 runs with different random seeds. The best results are marked in bold, and the second-best results are underlined. We calculate the $p$-values between the average performance of our method and the best performing baseline in both the metrics.
@@
[…]
$p$-value $<5e-2$ (Dice)
Table 6: Ablation study of the effects of different key components in our MAE + UNETR++ approach on the Synapse CT dataset. The reported values are the averages of 5 runs with different random seeds.
@@
UNETR++ 87.22 7.53
UNETR++ with pre-trained ViT encoder $\surd$ 87.62 7.03
UNETR++ with pre-trained ViT encoder & spatial pre-text task $\surd$ $\surd$ 88.09 6.12
UNETR++ with pre-trained ViT encoder & spatial pre-text task & topological loss $\surd$ $\surd$ $\surd$ 88.34 5.96
MAE + UNETR++ $\surd$ $\surd$ $\surd$ $\surd$ 88.94 5.89
4.3 Evaluation Metrics
We use the Dice score and HD95 as the evaluation metrics to measure the performances of the methods. For a specific semantic class, let $GT_{i}$ and $P_{i}$ denote the GT and prediction values for a voxel $i$, and $GT$ and $P$ denote the GT and prediction point sets, respectively.
The Dice score is defined as:
$\text{Dice}({GT},{P})=\frac{2\sum_{i}{GT_{i}}\times{P_{i}}}{\sum_{i}{GT_{i}}+\sum_{i}{P_{i}}}.$ (7)
The Hausdorff distance (HD) is defined as:
$\text{HD}({GT},{P})=\max\{\max_{g\in GT}\min_{p\in P}||g-p||,\max_{p\in P}\min_{g\in GT}||p-g||\}.$ (8)
The HD95 uses the 95th percentile of the distances between the GT and prediction point sets.
4.4 Experimental Results
@@
Synapse CT Dataset Results. In Table 1, we compare our method with an array of baseline methods (U-Net [50], TransUNet [56], UNETR [57], Swin-UNet [55], MISSFormer [58], Swin UNETR [59], and nnFormer [2]) and SOTA models (UNETR++ [3], and the MAE-based self pre-training method, i.e., UNETR + MAE [12]). On this dataset, UNETR++ yields superior performance over the other known methods. Our method outperforms UNETR++ by 1.72% and 1.64 mm in average Dice and HD95, respectively, which are quite impressive improvements on the Synapse CT dataset. Specifically, our method achieves the highest Dice scores on six organs (kidney (right), kidney (left), gallbladder, liver, stomach, and aorta). Compared to the known methods, our method is more advantageous in segmenting gallbladder, which is difficult to delineate using known segmentation methods. Moreover, our method is able to surpass the MAE-based self pre-training method, UNETR + MAE, by large margins in both the evaluation metrics, demonstrating the effectiveness of our method.
@@
BTCV CT Dataset Results. Table 2 showcases the segmentation results of various methods on the BTCV CT dataset. Among the known methods, nnU-Net [1] and UNETR++ [3] achieve average Dice scores of 83.16% and 83.28%, respectively. Our method outperforms the SOTA method UNETR++ by 0.8% in average Dice. This is particularly commendable given the challenging nature of the BTCV CT dataset, which encompasses 13 distinct organs.
@@
ACDC Dataset Results. Table 3 reports the quantitative results on the ACDC dataset. We observe that nnFormer [2] and UNETR++ [3] attain better performances of 92.06% and 92.83% in average Dice, respectively. Remarkably, our method surpasses the SOTA method UNETR++ by 0.62% in the average Dice score. Furthermore, our method outperforms the MAE-based self pre-training method UNETR + MAE [12] by an impressive 3.15% in average Dice, confirming the effectiveness of our new approach.
@@
MSD Spleen Dataset Results. As Table 4 shows, on the MSD spleen dataset, both UNETR + MAE [12] and UNETR++[3] already achieve a very high 0.966 Dice score, giving a limited margin for big improvement. Nonetheless, our method still manages to enhance both the Dice and HD95 scores by 0.8% and 0.244, respectively. These performances reinforce the superiority of our method over known SOTA methods.
@@
MSD Lung Dataset Results. Table 5 presents the experimental results on the MSD lung dataset. One can see that the best known method is UNETR++ [3], whose Dice score is higher than the second-best method, UNETR + MAE [12], by a margin of 1.78%. In comparison, our method outperforms UNETR++ by 1.87% and UNETR + MAE by a notable 3.65% in Dice score. These substantial improvements validate the effectiveness of our method.
Statistical Significance. In Tables 1, 2, 3, 4, and 5, we employ the independent two-sample $t$-test to calculate the $p$-values between the average performance of our method and the best-performing baseline in the evaluation metrics. The null hypothesis is that our method offers no superior benefits over the best-performing baseline. As one can see from these five tables, on all the five public datasets, our method consistently delivers $p$-values below $25e-$ 2, whether measured by HD95 and Dice or solely by Dice. This provides compelling evidence against the null hypothesis. Hence, our method yields significant improvements over known SOTA baseline methods on the five public datasets.
Fig. 3: Visual comparison of some qualitative segmentation results by our proposed method, UNETR++, UNETR + MAE, and UNETR on the Synapse CT dataset. The results of our method show its efficacy in more accurate segmentation. RKid: right kidney; Liv: liver; Aor: aorta; PSV: portal and splenic veins; Pan: pancreas; Sto: stomach; LKid: left kidney; IVC: inferior vena cava; Gal: gallbladder.
Fig. 4: Visual comparison of some qualitative segmentation results by our proposed method, UNETR++, nnFormer, UNETR + MAE, and UNETR on the ACDC dataset. The results highlight the high accuracy of our method. RV: right ventricle; LV: left ventricle; MYO: myocardium.
4.5 Ablation Study
@@
We conduct ablation study using the Synapse CT dataset to examine the effects of different key components in our method. From the results in Table 6, we observe the following. (1) When applying the pre-trained ViT encoder on top of UNETR++, the Dice is improved by $0.4\%$ ($p$-value $=0.027$, $t$-test), showing the effect of our fine-tuned model. (2) When adding the spatial pre-text task to the pre-training of ViT only, the Dice is further improved by $0.47\%$ ($p$-value $=0.009$, $t$-test), demonstrating the effect of our proposed pre-text task in extracting spatial information. (3) When further adding the topological loss to the pre-training of ViT only, the Dice is further improved by $0.25\%$ ($p$-value $=0.015$, $t$-test), validating the effect of our proposed topological loss in capturing geometric shape information. (4) When transferring the weights of both the ViT encoder and UNETR++ to the fine-tuned model, the Dice is improved by $0.60\%$ ($p$-value $=0.011$, $t$-test), validating the effects of the MAE pre-training strategy on UNETR++ as well as the spatial and reconstruction consistency losses on co-pretraining ViT and UNETR++.
4.6 Visualization of Segmentation Results
Fig. 3 displays some qualitative results by our method, UNETR++, UNETR + MAE, and UNETR on the Synapse CT dataset. Compared to the prediction results produced by the other methods, our prediction results are much closer to the ground truth (GT), showing clear advantages. Specifically, our method substantially reduces false positive predictions. This demonstrates the capability of our method in capturing geometric shape information and extracting spatial information.
Fig. 4 provides a qualitative comparison of segmentation results by UNETR++, nnFormer, UNETR + MAE, UNETR, and our method on the ACDC dataset. In the first row, one can see that the baseline methods, UNETR++, nnFormer, UNETR + MAE, and UNETR, generate some false positive predictions. However, our method does not. This validates that our method is able to capture geometric (i.e., topological) information with the proposed topological loss. In the second row, we observe that (1) UNETR and UNETR + MAE fail to segment the complete shape of the myocardium, (2) nnFormer produces some incorrect right ventricle pixels inside the myocardium, and (3) UNETR++ misses segmenting some myocardium pixels along the boundary. In contrast, our method generates highly accurate predictions, further showing its effectiveness.
5 Conclusions

### v3

@@ III-A Datasets and Experimental Setup
We conduct experiments on five segmentation datasets: Synapse multi-organ CT segmentation (Synapse CT Dataset) [15], BTCV multi-organ CT segmentation (BTCV CT Dataset) [15], ACDC automated cardiac diagnosis (ACDC) [3], and Medical Segmentation Decathlon (MSD) datasets [2] for two different segmentation tasks, spleen segmentation and lung segmentation. For each experiment, we perform 5 runs using different random seeds and report the average results. Additionally, we compute $p$-values to ascertain the statistical significance of the results.
@@
Synapse CT Dataset: This dataset [15] contains $30$ abdominal CT volumes with $8$ organs. Following [5, 24], we split the dataset randomly into $18$ volumes and $12$ volumes for training and testing, and report the average Dice and $95\%$ Hausdorff distance (HD95) on $8$ abdominal organs: spleen (Spl), right kidney (RKid), left kidney (LKid), gallbladder (Gal), liver (Liv), stomach (Sto), aorta (Aor), and pancreas (Pan).
@@
BTCV CT Dataset:: This dataset [15] consists of $30$ abdominal CT volumes with $13$ organs, including 8 organs of the Synapse CT dataset, along with esophagus (Eso), inferior vena cava (IVC), portal and splenic veins (PSV), right adrenal gland (RAG), and left adrenal gland (LAG).
@@
ACDC Dataset: This dataset [3] contains 100 samples, and aims to segment the cavity of the right ventricle, the myocardium of the left ventricle, and the cavity of the left ventricle. Each sample’s labels involve left ventricle (LV), right ventricle (RV), and myocardium (MYO). Following [27], we split the dataset into 70 training samples, 10 validation samples, and 20 test samples.
@@
MSD Spleen Dataset: This dataset [2] contains $41$ CT volumes for spleen segmentation. Following [10], we split the dataset into training, validation, and test sets (80:15:5).
@@
MSD Lung Dataset:: This dataset [2] comprises 64 CT volumes for lung cancer segmentation. We split the dataset with a 80:20 ratio for training and testing following [18].
III-B Implementation Details
@@
Our experiments are implemented with PyTorch and MONAI.The model training is performed on an NVIDIA Tesla V100 Graphics Card with 32GB GPU memory using the AdamW optimizer with a weight decay = $0.005$.
@@
For the Synapse CT and BTCV CT datasets, we clip the raw values between 175-175 and $250$, normalize the values into the range of [0, 1], and re-sample the spacing to $[1.5,1.5,2.0]$. All the models are trained with input images of size $96\times 96\times 96$. For the ACDC dataset, we re-sample the spacing to $[1.52,1.52,6.35]$. All the models are trained with input of size $160\times 160\times 16$. For the MSD spleen dataset, we clip the raw values between 57-57 and $164$, normalize the values into the range of [0, 1], and re-sample the spacing to $[1.5,1.5,2.0]$. All the models are trained with input of size $96\times 96\times 96$. For the MSD lung dataset, we clip the raw values between 1000-1000 and $3071$, normalize the values into the range of [0, 1], and re-sample the spacing to $[1.0,1.0,1.0]$. All the models are trained with input of size $192\times 192\times 32$.
@@
We use a learning rate of 6.4e36.4e-3 for self pre-training on all the datasets. We pre-train on the Synapse CT, BTCV CT, and MSD spleen and lung segmentation datasets with $10,000$ epochs, and on the ACDC dataset with $2,000$ epochs.
@@
For all the downstream segmentation tasks, we use a learning rate of 1e11e-1, and fine-tune with $5,000$ epochs for the Synapse CT, BTCV CT, and MSD spleen and lung segmentation datasets, and $1,000$ epochs for the ACDC dataset. The batch size for each case is set as the maximum size allowed by the GPU. We set $\lambda_{1}=0.5$, $\lambda_{2}=0.1$, and $\lambda_{3}=0.1$.
TABLE II: Segmentation results on the BTCV CT dataset.
@@
MAE + UNETR++ (ours) 94.97 87.93 87.37 78.46 78.97 96.99 88.31 92.51 89.01 76.94 80.18 69.88 71.48 84.08 $(0.8\%\uparrow)$
$p$-value <1e2<1e-2 (Dice)
TABLE III: Segmentation results on the ACDC dataset.
@@
$p$-value <1e2<1e-2 (Dice)
TABLE IV: Segmentation results on the MSD spleen dataset.
@@
MAE + UNETR++ (ours) 0.974 $(0.8\%\uparrow)$ 1.002 $(0.244\uparrow)$
$p$-values <1e2<1e-2 (Dice), <5e2<5e-2 (HD95)
TABLE V: Segmentation results on the MSD lung dataset.
@@
MAE + UNETR++ (ours) 82.55 $(1.87\%\uparrow)$
$p$-value <5e2<5e-2 (Dice)
TABLE VI: Ablation study of the effects of different key components in our MAE + UNETR++ approach on the Synapse CT dataset.
@@
UNETR++
87.22 7.53
UNETR++
with pre-trained ViT encoder
$\surd$ 87.62 7.03
UNETR++
with pre-trained ViT encoder
& spatial pre-text task
$\surd$ $\surd$ 88.09 6.12
UNETR++
with pre-trained ViT encoder
& spatial pre-text task
& topological loss
$\surd$ $\surd$ $\surd$ 88.34 5.96
MAE + UNETR++
$\surd$ $\surd$ $\surd$ $\surd$ 88.94 5.89
III-C Experimental Results
@@
Synapse CT Dataset Results. In Table I, we compare our method with an array of baseline methods (U-Net [17], TransUNet [5], UNETR [10], Swin-UNet [4], MISSFormer [12], Swin UNETR [9], and nnFormer [27]) and SOTA models (UNETR++ [18], and the MAE-based self pre-training method, i.e., UNETR + MAE [28]). On this dataset, UNETR++ yields superior performance over the other known methods. Our method outperforms UNETR++ by 1.72% and 1.64 mm in average Dice and HD95, respectively, which are quite impressive improvements on the Synapse CT dataset. Specifically, our method achieves the highest Dice scores on six organs (kidney (right), kidney (left), gallbladder, liver, stomach, and aorta). Compared to the known methods, our method is more advantageous in segmenting gallbladder, which is difficult to delineate using known segmentation methods. Our method is able to surpass the UNETR + MAE by large margins in both the evaluation metrics, demonstrating the effectiveness of our method (see Fig. 3).
@@
BTCV CT Dataset Results. Table II showcases the segmentation results of various methods on the BTCV CT dataset. Among the known methods, nnU-Net [13] and UNETR++ [18] achieve average Dice scores of 83.16% and 83.28%, respectively. Our method outperforms the SOTA method UNETR++ by 0.8% in average Dice. This is particularly commendable given the challenging nature of the BTCV CT dataset, which encompasses 13 distinct organs.
@@
ACDC Dataset Results. Table III reports the quantitative results on the ACDC dataset. We observe that nnFormer [27] and UNETR++ [18] attain better performances of 92.06% and 92.83% in average Dice, respectively. Remarkably, our method surpasses the SOTA method UNETR++ by 0.62% in the average Dice score. Furthermore, our method outperforms the MAE-based self pre-training method UNETR + MAE [28] by an impressive 3.15% in average Dice, confirming the effectiveness of our new approach.
@@
MSD Spleen Dataset Results. As Table IV shows, on the MSD spleen dataset, both UNETR + MAE [28] and UNETR++[18] already achieve a very high 0.966 Dice score, giving a limited margin for big improvement. Nonetheless, our method still manages to enhance both the Dice and HD95 scores by 0.8% and 0.244, respectively. These results reinforce the superiority of our method over known SOTA methods.
@@
MSD Lung Dataset Results. Table V presents the experimental results on the MSD lung dataset. One can see that the best known method is UNETR++ [18], whose Dice score is higher than the second-best method, UNETR + MAE [28], by a margin of 1.78%. In comparison, our method outperforms UNETR++ by 1.87% and UNETR + MAE by a notable 3.65% in Dice score. These substantial improvements validate the effectiveness of our method.
Figure 3: Visual results of different methods on the Synapse CT dataset.
III-D Ablation Study
@@
We conduct ablation study using the Synapse CT dataset to examine the effects of different key components in our method. From the results in Table VI, we observe the following. (1) When applying the pre-trained ViT encoder on top of UNETR++, the Dice is improved by $0.4\%$ ($p$-value =0.027=0.027, $t$-test), showing the effect of our fine-tuned model. (2) When adding the spatial pre-text task to the pre-training of ViT only, the Dice is further improved by $0.47\%$ ($p$-value =0.009=0.009, $t$-test), demonstrating the effect of our proposed pre-text task in extracting spatial information. (3) When further adding the topological loss to the pre-training of ViT only, the Dice is further improved by $0.25\%$ ($p$-value =0.015=0.015, $t$-test), validating the effect of our proposed topological loss in capturing geometric shape information. (4) When transferring the weights of both the ViT encoder and UNETR++ to the fine-tuned model, the Dice is improved by $0.60\%$ ($p$-value =0.011=0.011, $t$-test), validating the effects of the MAE pre-training strategy on UNETR++ as well as the spatial and reconstruction consistency losses on co-pretraining ViT and UNETR++.
IV Conclusions
@@
V Acknowledgements
This research was supported in part by NSF Grants CCF-2523787 and OAC-2104158.